OpenStudy (anonymous):

What is the arc length of a circle that has a 6-inch radius and a central angle that is 65 degrees? Use 3.14 for π and round your answer to the nearest hundredth. - 0.65 inch - 1.13 inches - 6.80 inches - 390.01 inches

OpenStudy (anonymous):

@Mertsj Can you help me please or ask anyone to?

OpenStudy (anonymous):

@wolf1728

OpenStudy (mertsj):

65 degrees is 65/360 of the entire circle. The length of the entire circle is 2 pi r. Find 2 pi r and multiply it by 65/360

OpenStudy (wolf1728):

6-inch radius central angle = 65 degrees Circumference = 2*PI*6 Circumference = 37.70 The arc would be (65/360) of the circle or 0.18055556 0.18055556 * 6 = 1.083333 Rounding pi to 3.14 would make the answer 1.08

OpenStudy (anonymous):

The closest answer would 1.13 because it is closer

OpenStudy (anonymous):

Can you help with another

OpenStudy (wolf1728):

Wait I messed up answer is 0.18055556 * 37.70 = 6.806784 SORRY !!!

OpenStudy (anonymous):

oh thank you lol

OpenStudy (anonymous):

can you help with another

OpenStudy (wolf1728):

Here's a calcualtor to check it http://www.1728.org/radians.htm And I'll help with another

OpenStudy (anonymous):

The point (−3, 1) is on the terminal side of angle θ, in standard position. What are the values of sine, cosine, and tangent of θ? Make sure to show all work.

OpenStudy (wolf1728):

I don't exactly know know specifically what they mean by terminal side, standard position. MAYBE it's that angle starts at (0,0)?

OpenStudy (anonymous):

Hmmm.....

OpenStudy (mertsj):